OpenStudy (anonymous):

How do I start to solve the triangle using Law of Cosines and information A=48 deg, b=3, and c=14 ? Thanks!

OpenStudy (anonymous):

\[a^2=b^2+c^2-2ab\cos(A)\]

OpenStudy (anonymous):

\[a^2=3^2+14^2-2\times 3\times 14\cos(48)\] and a calculator will find \(c^2\) take the square root to find \(c\)

OpenStudy (anonymous):

here is \(c^2\) http://www.wolframalpha.com/input/?i=3^2%2B14^2-2*3*14cos%2848%29